OpenStudy (anonymous):

The height of an arrow shot into the air is nodeled by the equation y=64t-16t^2, where y=feet above the ground and t seconds after release. How highdoes the arrow go?

OpenStudy (anonymous):

32 metres

OpenStudy (anonymous):

Show work?

OpenStudy (anonymous):

on differertitating u will get the ans

OpenStudy (anonymous):

I am only in algebra I

OpenStudy (anonymous):

@dldinatale wat can i do thn..? u want me to do it by algebra...?

OpenStudy (anonymous):

Yes. We. eare on quadratic equations in 9th grade

OpenStudy (anonymous):

take 16 common frm right side nw u will get 16[4t-t^2]=y

OpenStudy (anonymous):

Just take the derivative. The derivative is just the rate of change of the equation when the instantaneous rate of change is zero the ball has stopped and will start falling down again. Just do nx^n-1 so here 64t^(1-1)-16*2t^(2-1) = 64-32t and then solve for when the equation equals zero and isolate t. This will give you the time when the ball is at it's max height then sub that time t back into the first equation which gave you the height at each moment and you will have the height right before the arrow starts to fall back down which is the max height. I got 64 but I might be doing something wrong does this look right to you nasavivek?